in parallelograms ABCD abovewhat is the value of x?
10x (25x-30)
: in parallelograms ABCD abovewhat is the value of x?
10x (25x-30)

Answer by jojo14344(1030)

10x (25x-30)
Distribute:
250x^2-300x=0 --------------------> working eqn
Divdide the whole eqn by 250;
cross(25)x^2/cross(250)-cross(300)1.2x/cross(250)=0/cross(250)
x^2-1.2x=0
x(x-1.2)=0
x=1.2
Check it out via working eqn:
250(1.2)^2-300(1.2)=0
250(1.44)-360=0
360-360=0
0=0
Thank you,
Jojo